# Contamination Detection for VLMs Using Multi‑Modal Semantic Perturbations

### v1

@@ 6 Analysis of Multi-modal Semantic Perturbation
Why perturbations yield a generalized benchmark. Our core assumption is that by preserving the original question and only altering the answer choice, the question difficulty remains comparable. In addition, we often observe that the perturbation highlights salient visual cues more clearly than the original images (Fig. 3), collectively yielding an alternate benchmark that is similar or easier. We validate this empirically: clean models consistently achieve higher accuracy on perturbed benchmarks (Table 2).
@@
Providing the question-answer pair in caption generation is critical. As described in Sec. 4, conditioning the caption generation on both the original question and the new answer is critical to creating a generalized benchmark. One natural approach is to simply create a variant of the original image without changing the answer to the question. To test this, we create a dense caption of the original image without providing any context and reconstruct the images. We empirically observe that generative models often fail to extract relevant information from the dense details of the original image, leading to generations that are equivalent in difficulty or inconsistent. Changing the answer is necessary to delineate contaminated models’ behavior from simply memorizing the answer.
Failure modes of multi-modal semantic perturbation. There are cases where multi-modal perturbation fails to reveal contamination. The perturbed image may differ in its visual details that it no longer closely resembles the original. In such cases, contaminated models may answer both the original and perturbed questions correctly as shown in Figure 4, hiding the model’s contamination.
@@
Both RealWorldQA and MMStar are multiple-choice benchmarks. Although the multiple-choice format simplifies evaluation, our framework is not inherently tied to this setting. Once visual evidence is perturbed, evaluation can be adapted to free-form tasks using string matching, likelihood-based scoring, or LLM-as-a-judge approaches. This principle extends naturally beyond multi-choice VQA.
Finally, manual filtering is required only because of current limitations of diffusion models. As generative models improve, we expect manual filtering to become unnecessary, further strengthening the scalability of our approach.
@@ 7 Ablation Studies
We next validate that the core idea—testing for memorization via mutli-modal semantic perturbations—is robust to design choices. Specifically, we show our method is not tied to synthetic edits, scales to larger models and alternate contamination regimes, and works without GPT-4o or manual filtering.
Real-world Counterfactuals: NaturalBench. NaturalBench (naturalbench) provides real counterfactual pairs – photos of the same scene under altered conditions – serving as a natural analogue to our synthetic variants. We fine-tune on one variant and evaluate its paired counterfactual with the same question. Contaminated models drop sharply (up to 45.95%), while clean models remain comparatively stable (Table 5), showing our detector generalizes beyond synthetic perturbations [^2] Note that because counterfactuals are not guaranteed to be easier, clean-model deltas need not be positive.. Thus, any reliable semantic variation – natural, procedural, or synthetic – fits our framework. Full results are listed in E.2.
Model Scale: LLaVA-13B. To assess scalability, we repeat the RealWorldQA experiment with LLaVA-v1.5-13B. Table 5 shows our detector remains effective at this scale. Given that larger models are more prone to memorization, these findings indicate the approach is suitable for stronger VLMs. Full results are listed in E.2.
@@
Table 4: Performance of clean and contaminated models on NaturalBench. While clean model shows comparable performance on the test set, contaminated models fail to generalize, demonstrating dramatic performance drop upto -45.58%
@@
Table 5: Performance of clean and contaminated LLaVA-v1.5-13B on RealWorldQA after multi-modal semantic perturbation. The result satisfies all three requirements.
@@
Finally, we ablate two key components of our pipeline to show modularity.
Automation of filtering Process. We replace manual validation with the o3 model to filter generated image–question pairs. The automatic pass removes 471 items and retains 294, of which 253 overlap with the manually kept set, indicating high agreement (Table 7). The exact prompt is provided in Appendix A.3 and full results are listed in E.3.
Caption generation with Molmo-7B-D. To decouple the pipeline from GPT-4o, we substitute captioning with the lightweight open-source Molmo-7B-D (molmo) while keeping Flux ControlNet for image generation. After manual filtering, this variant yields 398 valid pairs and preserves the same detection trends, underscoring the flexibility of our approach (Table 8). Full results are listed in E.4.
@@
LLM+MLP (contaminated) 1 58.50 54.42 –4.08
2 60.54 54.42 –6.12
3 64.97 53.40 –11.57
Table 7: Performance after filtering with the o3 model, resulting in 294 valid image–question pairs. Results satisfy all requirements.
@@
Table 8: Performance after generating and filtering images with captions from Molmo-7B-D, resulting in 398 valid image–question pairs. Results satisfy all requirements.
@@ 8 Related Work
Data Contamination in LLMsgpt3 was among the first to highlight the problem of data contamination when training models on internet-scale corpora, proposing an n-gram overlap-based decontamination technique that was later adopted in abdin2024phi4technicalreport. Building on this, several detection methods exploit verbatim memorization to identify leaked test-set examples (provingtestset; guided-prompting; ngram-acc). More specifically, ngram-acc; guided-prompting test whether the model can accurately reconstruct masked spans of test questions, while provingtestset measure the log-probability of the original ordering of multiple-choice options. choi-conta instead examine embedding divergence after fine-tuning, exploiting the observation that embeddings of unseen samples change more substantially than those of memorized ones. However, these methods face key limitations: provingtestset; guided-prompting; ngram-acc perform poorly on contaminated VLMs, while choi-conta requires ground-truth clean model behavior to establish a detection threshold, reducing its practicality.
Generalized Benchmarks for Detecting Contaminationyao-dcclb address these limitations by testing for generalization rather than memorization. They construct trivial variants of benchmark questions by replacing incorrect multiple-choice options with irrelevant answers. Contaminated models often fail on the easier variants, exposing strong memorization. However, their setup involves training for 36 epochs on a single benchmark – an unrealistic scenario for modern large-scale training. In a similar spirit, GSM-Symbolic and MATH-Perturb (gsm-symbolic; mathperturb) introduce perturbation-based approaches, measuring performance drops on variant questions as contamination signals.
Contamination Detection in VLMs. VLMs differ from LLMs due to their multi-modal inputs and multi-stage alignment training, which yield distinct contamination dynamics. vlm-cleaneval proposed shuffling BGR color channels to mitigate spurious cues, while vlm-systematic introduced image masking and option shuffling to test robustness. Yet these approaches struggle to reliably detect contaminated VLMs, as observed performance drops may arise from confounding factors such as visual artifacts or biased sampling rather than true memorization. In contrast, our framework provides consistent detection across diverse fine-tuning regimes, requires no access to leaked data, and yields performance drops that correlate strongly with the degree of contamination.
@@ 9 Conclusion
Reproducibility Statement. We will publicly release our code, models and data.
@@ A.2 Multi-modal Semantic Perturbation
A.3 System Prompts to GPT-4o and o3
To generate the semantically perturbed question, we utilize GPT-4o gpt4o and Flux diffusion model flux2024 trained on ControlNet Zhang2023ConditionalControl. We generate a detailed caption about the image using GPT-4o along with the original question and the answer with the following system prompt:
@@ A.3 System Prompts to GPT-4o and o3
Appendix B Perplexity-based Methods
We evaluate three perplexity-based methods: Shared Likelihood (SL) (provingtestset), which compares log-likelihoods on original versus shuffled inputs; Guided Prompting (GP) (guided-prompting), which uses an LLM to score model completions against the original; and N-gram Accuracy (NG) (ngram-acc), which masks answer choices and checks reproduction. Each probes whether models memorize test samples through partial or perturbed inputs in the text domain.
@@ Appendix B Perplexity-based Methods
Appendix C Detecting Contaminated Models with Multi-Modal Semantic Perturbation on RealWorldQA
@@
Table 11: Performance of LLaVA-v1.5-7B (top) and Qwen2-VL-7B (bottom) on the RealWorldQA dataset. We compare to “Multi‐modal Leakage” mmstar, CircularEval (mmbench), and “Choice Confusion” yao-dcclb. Clean models perform better – confirming that the perturbed questions are indeed of equal or lower difficulty. In contrast, all contaminated models perform worse, enabling reliable detection by our method via a simple check for performance drops. RQA denotes RealWorldQA and "_P" denotes the semantically perturbed version; "to" denotes text-only performance; "_C" denotes evaluation using circular options; "_G" denotes evaluation using choice confusion; $\Delta$ denotes the difference in performance, with positive values indicating gains. "Success?" indicates whether the method detected contamination. "Require clean model?" indicates whether the method requires access to a clean model as a baseline. If a clean model is required as reference, the method cannot be used to detect the reference models themselves, so the entry is marked as “–”.
Appendix D Extended Evaluation on Additional Open-Source and Proprietary Models
We further apply our pipeline to GPT-4o (gpt4o), Gemini-2.0-Flash (geminiteam2024gemini), Phi-3.5-V (abdin2024phi), and InternVL-2.5 (internvl2.5). We do not contaminate these models. Although we cannot guarantee that these models have never encountered our evaluation data, we assume they are uncontaminated. Under this assumption, if our framework is sound, their performance should remain consistent across original and perturbed datasets. Indeed, our results in Table 12 confirm this expectation. Across all models, perturbed performance exceeds original performance, reinforcing that clean models generalize while contaminated ones fail. Notably, Phi-3.5-V exhibits the largest gain, while InternVL-2.5-8B remains relatively flat, demonstrating that our metric is robust across architectures. This consistency underscores the effectiveness of multi-modal semantic perturbation as a detection framework.
@@ Appendix D Extended Evaluation on Additional Open-Source and Proprietary Models
Appendix E Full Experiment Results
@@
E.1 Full Results on LLaVA-v1.5-7B  and Qwen2-VL-7B
@@
Table 141416, and 16 list out all results for four varying training strategies for LLaVA-v1.5-7B  and Qwen2-VL-7B , including the non-default training strategies for each model which were omitted due to limited space. Note that the results here are computed on 440 manually filtered images. Our approach detects contaminated models regardless of the training strategies and satisfies all three requirements.
E.2 Full Results with LLaVA-v1.5-13B and NaturalBench
Table 18 lists full results for LLaVA-v1.5-13B model trained on RealWorldQA. Note that the results here are computed on 440 manually filtered images. Table 18 lists full results for LLaVA-v1.5-7B  traiend on one variant of NaturalBench and tested on the counterfactual version. The clean and consistent detection of contaminated models show that our approach can be applied to models of larger scale and to real-world perturbations.
@@
E.3 Full Results with o3 filtering

### v2

@@ 6 Analysis of Multi-modal Semantic Perturbation
Why perturbations yield a generalized benchmark. Our core assumption is that by preserving the original question and only altering the answer choice, the question difficulty remains comparable, and that clean models that answer the original question correctly will solve the variant. In addition, we often observe that the perturbation highlights salient visual cues more clearly than the original images (Fig. 3), collectively yielding an alternate benchmark that is similar or easier. We validate this empirically: clean models consistently achieve higher accuracy on perturbed benchmarks (Table 2).
@@
Providing the question-answer pair in caption generation is critical. As described in Sec. 4, conditioning the caption generation on both the original question, original answer and the new answer is critical to creating a generalized benchmark. One natural approach is to create a variant of the original image from simply conditioning the generated caption on the question and the new correct answer. When we evaluated clean models on images generated from this version of the prompt, the clean model performance was much lower compared to both the original dataset and our final perturbed dataset, with invalid perturbations appearing much more frequently, for example, due to critical component of the image being left out so the question is ambiguous or no longer solvable. Intuitively, the captioning model first reasons about which parts of the image need to change to make the new answer correct, and produces a detailed caption emphasizing those changes. Flux+ControlNet then focuses on rendering these components exactly. Finally, changing the answer is necessary to delineate contaminated models’ behavior from simply memorizing the answer.
Failure modes of multi-modal semantic perturbation. There are cases where multi-modal perturbation fails to reveal contamination. The perturbed image may differ in its visual details that it no longer closely resembles the original. In such cases, contaminated models may answer both the original and perturbed questions correctly as shown in Figure 4, hiding the contamination of the model. To ensure that such failure cases are rare, we manually inspected the perturbed datasets and verified that only 8 out of 440 images ($\sim$1.8%) from perturbed RealWorldQA and 17 out of 495 images ($\sim$3.4%) from perturbed MMStar deviate from the original question’s visual details.
@@
Both RealWorldQA and MMStar are multiple-choice benchmarks. Although the multiple-choice format simplifies evaluation, our framework is not inherently tied to this setting. Once visual evidence is perturbed, evaluation can be adapted to free-form tasks using string matching, likelihood-based scoring, or LLM-as-a-judge approaches. This principle extends naturally beyond multi-choice VQA, such as free-form QA settings, as shown in Appendix F.7.
Finally, manual filtering is required only because of current limitations of diffusion models. As generative models improve, we expect manual filtering to become unnecessary.
@@ 7 Ablation Studies
We next validate that the core idea – testing for memorization via mutli-modal semantic perturbations – is robust to design choices. Specifically, we show our method is not tied to synthetic edits, generalizes to paraphrased data, scales to larger models and alternate contamination regimes, and works without GPT-4o or manual filtering.
Real-world counterfactuals. NaturalBench (Li et al., 2024) provides real counterfactual pairs – photos of the same scene under altered conditions – serving as a natural analogue to our synthetic variants. We fine-tune on one variant and evaluate on its paired counterfactual. Contaminated models drop sharply (up to 45.95%), while clean models remain comparatively stable (Table 7), showing our method generalizes beyond synthetic perturbations [^3] Note that because counterfactuals are not guaranteed to be easier, clean-model deltas need not be positive.. Thus, any reliable semantic variation – natural, procedural, or synthetic – fits our framework. Full results are listed in F.2.
Contamination with paraphrased data. A simple but effective contamination is paraphrasing the data the model is contaminated on. To test whether models contaminated with paraphrased data can be detected, we paraphrase the questions using GPT-4o with minimal n-gram overlap. The contaminated models are then evaluated on the original benchmark and its perturbed variant. Table 4 and 5 show that contaminated models show inflated performance on original benchmarks with lowered performance on perturbed benchmarks, proving our detection method is robust against paraphrasing.
Method Epoch RQA RQA_P $\Delta$
Contamination with paraphrased RealWorldQA
LLaVA-v1.5-7B (clean) 52.05 56.36 +4.31
LoRA (contaminated) 1 52.03 40.00 -12.03
2 55.91 40.00 -15.91
3 59.09 38.86 -20.23
LLM+MLP (contaminated) 1 56.14 53.41 -2.73
2 61.36 48.64 -12.72
3 63.64 49.77 -13.87
Table 4: Performance of LLaVA-v1.5-7B and contaminated variants on RealWorldQA when trained on the paraphrased benchmark. “_P” denotes the semantically perturbed version. For full results, refer to Appendix F.5.
Method Epoch MMStar MMStar_P $\Delta$
Contamination with paraphrased MMStar
LLaVA-v1.5-7B (clean) 37.78 68.29 +31.51
LoRA (contaminated) 1 51.52 43.64 -7.88
2 59.60 41.82 -17.78
3 62.22 40.81 -21.41
LLM+MLP (contaminated) 1 47.27 46.06 -1.13
2 57.60 53.54 -3.06
3 62.42 57.98 -4.44
Table 5: Performance of LLaVA-v1.5-7B and contaminated variants on MMStar when trained on the paraphrased benchmark. “_P” denotes the semantically perturbed version. For full results, refer to Appendix F.5.
Model scale: LLaVA-13B. To assess scalability, we repeat the RealWorldQA experiment with LLaVA-v1.5-13B. Table 7 shows our detector remains effective at this scale. Given that larger models are more prone to memorization, these findings indicate the approach is suitable for stronger VLMs. Full results are listed in F.2.
@@
Table 6: Performance of clean and contaminated models on NaturalBench. While clean model shows comparable performance on the test set, contaminated models fail to generalize, with performance drop upto -45.58%. For full results, refer to Appendix F.2.
@@
Table 7: Performance of clean and contaminated LLaVA-v1.5-13B on RealWorldQA after multi-modal semantic perturbation. The result satisfies all three requirements. For full results, refer to Appendix F.2.
@@
Test-set leakage with a mixture of benchmarks We test our perturbation pipeline under a more realistic contamination scenario: a mixture of other benchmarks. We fine-tune the models with popular multi-modal benchmarks: MathVista Lu et al. (2024b), MMMU Yue et al. (2024), MMBench Liu et al. (2024b) and CV-Bench Tong et al. (2024), resulting in 11,280 image-question pairs, with RealWorldQA and MMStar now consisting only of $\sim$6.7% and $\sim$13.3%, respectively. We still observe a clean and consistent detection across all contaminated models, demonstrating that our perturbation pipeline can reliably detect contaminated models even when the contamination signal is weaker. Results are in Appendix F.6 due to limited space.
Finally, we ablate two key components of our pipeline to show modularity of our pipeline.
Automation of filtering process. We replace manual validation with the o3 model to filter generated image–question pairs. The automatic pass removes 471 items and retains 294, of which 253 overlap with the manually kept set, indicating high agreement (Table 9). The exact prompt is provided in Appendix B and full results are listed in F.3.
Caption generation with Molmo-7B-D. To decouple the pipeline from GPT-4o, we substitute captioning with the lightweight open-source Molmo-7B-D (Deitke et al., 2024). After manual filtering, this variant yields 398 valid pairs and preserves the same detection trends, underscoring the flexibility of our approach (Table 10). Full results are listed in F.4.
@@
LLM+MLP (contaminated) 1 60.54 54.42 –6.12
2 63.95 55.10 –8.85
3 64.63 53.40 –11.23
Table 9: Performance after filtering with the o3 model, resulting in 294 valid image–question pairs. Results satisfy all requirements. For full results, refer to Appendix F.3.
@@
Table 10: Performance after generating and filtering images with captions from Molmo-7B-D, resulting in 398 valid image–question pairs. Results satisfy all requirements. For full results, refer to Appendix F.4.
@@ 8 Related Work
Data contamination in LLMs. Brown et al. (2020) was among the first to highlight the problem of data contamination when training models on internet-scale corpora, proposing an n-gram overlap–based decontamination technique that was later adopted in Bai et al. (2023). Building on this, several detection methods exploit verbatim memorization to identify leaked test-set examples (Oren et al., 2023; Golchin and Surdeanu, 2024; Xu et al., 2024). More specifically, Xu et al. (2024); Golchin and Surdeanu (2024) test whether the model can accurately reconstruct masked spans of test questions, while Oren et al. (2023) measure the log-probability of the original ordering of multiple-choice options. Choi et al. (2025) instead examine embedding divergence after fine-tuning, exploiting the observation that embeddings of unseen samples change more substantially than those of memorized ones. However, these methods face key limitations: Oren et al. (2023); Golchin and Surdeanu (2024); Xu et al. (2024) perform poorly on contaminated VLMs, while Choi et al. (2025) requires ground-truth clean model behavior to establish a detection threshold, reducing its practicality.
Generalized benchmarks for detecting contamination. Yao et al. (2024) tests for generalization rather than memorization. They construct trivial variants of benchmark questions by replacing incorrect multiple-choice options with irrelevant answers. Contaminated models often fail on the easier variants, exposing strong memorization. However, their setup involves training for 36 epochs on a single benchmark – an unrealistic scenario for modern large-scale training. Similarly, GSM-Symbolic and MATH-Perturb (Mirzadeh et al., 2024; Huang et al., 2025) introduce perturbation-based approaches, measuring performance drops on variant questions as contamination signals.
Contamination detection in VLMs. VLMs differ from LLMs due to their multi-modal inputs and multi-stage alignment training, which yield distinct contamination dynamics. Lu et al. (2024a) proposed shuffling BGR color channels to mitigate spurious cues, while Song et al. (2025) introduced image masking and option shuffling to test robustness. Yet these approaches struggle to reliably detect contaminated VLMs, as observed performance drops may arise from confounding factors such as visual artifacts or biased sampling rather than true memorization. In contrast, our framework provides consistent detection across diverse fine-tuning regimes, requires no access to leaked data, and yields performance drops that correlate strongly with the degree of contamination.
@@ 9 Conclusion
Reproducibility statement. We publicly release our code, models and data.
LLM usage. Gemini and ChatGPT were used to polish some of the paper’s writing.
@@ A.2 Multi-modal Semantic Perturbation
A.3 Cost of Multi-modal semantic perturbation
First, we note that our semantic perturbation pipeline only needs to be run once to constructed the perturbed benchmark. We also note that the cost scales linearly with the size of the dataset. We list below the cost and hardware requirements of each stage of our pipeline.
Caption generation. Using GPT-4o as the captioning model required on average $\sim$75 input tokens and $\sim$125 output tokens per image. This means that generating 1,000 captions would cost less than \$1.50 (USD). This stage also does not require proprietary models: it can be replaced with Molmo-7B-D (Table 102424), which fits on a single 24GB GPU.
Flux+ControlNet. Flux+ControlNet generation runs on a single 40GB GPU. With 25 sampling steps, each image takes around 15 20 seconds to generate, and this can be further reduced by decreasing the number of steps. This stage is trivially parallelizable across multiple GPUs or machines if higher throughput is needed.
Automated filtering with o3. Automated filtering required on average $\sim$75 input tokens and $\sim$200 output tokens per image. Generating 1,000 filtering decisions costs less than \$2.00 (USD).
Overall, the total cost to construct the perturbed benchmark is cheap with modest hardware requirements. Our approach scales linearly with dataset size, making our approach practical and scalable even for reasonably large benchmarks. Finally, when real-world counterfactuals are already available as part of the benchmark design (e.g., NaturalBench in Table 720), we do not need to generate perturbations at all.
Appendix B System Prompts to GPT-4o and o3
To generate the semantically perturbed question, we utilize GPT-4o OpenAI (2024) and Flux diffusion model Labs (2024) trained on ControlNet Zhang et al. (2023). We generate a detailed caption about the image using GPT-4o along with the original question, original answer and the new correct answer with the following system prompt:
@@ Appendix B System Prompts to GPT-4o and o3
Appendix C Perplexity-based Methods
We evaluate three perplexity-based methods: Shared Likelihood (SL) (Oren et al., 2023), which compares log-likelihoods on original versus shuffled inputs; Guided Prompting (GP) (Golchin and Surdeanu, 2024), which uses an LLM to score model completions against the original; and N-gram Accuracy (NG) (Xu et al., 2024), which masks answer choices and checks reproduction. Each probes whether models memorize test samples through partial or perturbed inputs in the text domain.
@@ Appendix C Perplexity-based Methods
Appendix D Detecting Contaminated Models with Multi-Modal Semantic Perturbation on RealWorldQA
@@
Table 13: Performance of LLaVA-v1.5-7B (top) and Qwen2-VL-7B (bottom) on the RealWorldQA dataset. We compare to “Multi‐modal Leakage” Chen et al. (2024a), CircularEval (Liu et al., 2024b), and “Choice Confusion” Yao et al. (2024). Clean models perform better – confirming that the perturbed questions are indeed of equal or lower difficulty. In contrast, all contaminated models perform worse, enabling reliable detection by our method via a simple check for performance drops. RQA denotes RealWorldQA and "_P" denotes the semantically perturbed version; "to" denotes text-only performance; "_C" denotes evaluation using circular options; "_G" denotes evaluation using choice confusion; $\Delta$ denotes the difference in performance, with positive values indicating gains. "Success?" indicates whether the method detected contamination. "Require clean model?" indicates whether the method requires access to a clean model as a baseline. If a clean model is required as reference, the method cannot be used to detect the reference models themselves, so the entry is marked as “–”. For full results, refer to Appendix F.
Appendix E Extended Evaluation on Additional Open-Source and Proprietary Models
We further apply our pipeline to GPT-4o (OpenAI, 2024), Gemini-2.0-Flash (Gemini Team, 2024), Phi-3.5-V (Abdin et al., 2024b), and InternVL-2.5 (Chen et al., 2025). We do not contaminate these models. Although we cannot guarantee that these models have never encountered our evaluation data, we assume they are uncontaminated. Under this assumption, if our framework is sound, their performance should remain consistent across original and perturbed datasets. Indeed, our results in Table 14 confirm this expectation. Across all models, perturbed performance exceeds original performance, reinforcing that clean models generalize while contaminated ones fail. Notably, Phi-3.5-V exhibits the largest gain, while InternVL-2.5-8B remains relatively flat, demonstrating that our metric is robust across architectures. This consistency underscores the effectiveness of multi-modal semantic perturbation as a detection framework.
@@ Appendix E Extended Evaluation on Additional Open-Source and Proprietary Models
Appendix F Full Experiment Results
@@
F.1 Full Results on LLaVA-v1.5-7B  and Qwen2-VL-7B
@@
Table 161618, and 18 list out all results for four varying training strategies for LLaVA-v1.5-7B  and Qwen2-VL-7B , including the non-default training strategies for each model which were omitted due to limited space. Note that the results here are computed on 440 manually filtered images. Our approach detects contaminated models regardless of the training strategies and satisfies all three requirements.
F.2 Full Results with LLaVA-v1.5-13B and NaturalBench
Table 20 lists full results for LLaVA-v1.5-13B model trained on RealWorldQA. Note that the results here are computed on 440 manually filtered images. Table 20 lists full results for LLaVA-v1.5-7B  trained on one variant of NaturalBench and tested on the counterfactual version. The clean and consistent detection of contaminated models show that our approach can be applied to models of larger scale and to real-world perturbations.
@@
F.3 Full Results with o3 filtering